# Spherical orbits around Kerr-Newman and Ghosh black holes

### v4

@@ Abstract
We conduct a comprehensive study on spherical orbits around two types of black holes: Kerr-Newman black holes, which are charged, and Ghosh black holes, which are nonsingular. In this work, we consider both null and timelike cases of orbits. Utilizing the Mino formalism, all analytical solutions for the geodesics governing these orbits can be obtained. It turns out that all spherical photon orbits outside the black hole horizons are unstable. In the extremal cases of both models, we obtain the photon boomerangs. The existence of charge in the Kerr-Newman allows the orbits to transition between retrograde and prograde motions, and its increase tends to force the orbits to be more equatorial. On the other hand, the Ghosh black hole, characterized by a regular core and a lack of horizons in certain conditions, presents the possibility of observable stable spherical orbits in the so-called no-horizon condition. As the Ghosh parameter $k$ increases, trajectories tend to exhibit larger latitudinal oscillation amplitudes. We observe that as the Ghosh parameter $k$ increases the trajectories tend to have larger latitudinal oscillation amplitudes. Finally, we investigate the existence of innermost stable spherical orbits (ISSOs). Both black holes demonstrate the appearance of two branches of ISSO radii as a function of the Carter constant C. However, there are notable differences in their behavior: in the case of the Kerr-Newman black hole, the branches merge at a critical value, beyond which no ISSO exists, while for the Ghosh black hole, the transcendental nature of the metric function causes the branches to become complex at some finite distance.
@@ I Introduction
It is well-known that any system possessing spherical symmetry and staticity will preserve its angular momentum, and as a result, the corresponding dynamics are confined on a plane. This is true in classical mechanics as well as in general relativity (GR). What is genuine in GR, among other things, is that not only massive particles but light can also orbit around a massive object. Perhaps the ultimate massive object in the universe is a black hole. Its visual image was first released in 2019 by the Event Horizon Telescope (EHT) collaboration EventHorizonTelescope:2019dse for M87*, and for Sgr A* three years later EventHorizonTelescope:2022wkp . For Schwarzschild we know that the photon circular orbit (photon sphere, $r_{ps}$) is $r_{ps}=3M$ Carroll:2004st , while for Reissner-Nordstrom (RN) there are $r_{ps}^{\pm}=3/2\left(M\pm\sqrt{M^{2}-8Q^{2}/9}\right)$ Khoo:2016xqv ; Pradhan:2010ws . Both the Schwarzschild and RN unstable photon spheres are outside their corresponding shadow radii Bardeen0 ; Luminet:1979nyg ; Falcke:1999pj ; Johannsen:2010ru ; Kumar:2018ple ; Afrin:2021wlj , thus can be observed. In reality, however, all astrophysical objects including black holes rotate. The appropriate metric solutions are the Kerr and the Kerr-Newman (KN) Kerr:1963ud ; Newman:1965my . EHT observed the Sgr A* angular diameter $d_{sh}=48.7\pm 7\mu as$, which exhibits consistency with theoretical predictions of Kerr black hole EventHorizonTelescope:2022wkp . Moreover, the observation confirms that the Sgr A* shows the viability of the Kerr model with constrained charge (KN) Ghosh:2022kit ; Vagnozzi:2022moj ; Khodadi:2022pqh ; KumarWalia:2022aop ; KumarWalia:2022ddq ; Uniyal:2022vdu .
@@
The exploration of ”spherical photon orbits” (SPO) began with Stuchlik’s investigation of the Kerr model Stuchlik1981 , followed by Calvani, De Felice, and Nobili’s study of the Kerr-Newman scenario Calvani:1980is . However, all studies on SPOs suggest that orbits are stable only within the inner horizon; beyond this region, they become inherently unstable. Nonetheless, SPOs play a crucial role in the observation of black holes. It is the multiple orbits of photon rays around black holes, terminated by their instability, that contribute to the formation of the luminous ring detected by instruments like the Event Horizon Telescope (EHT) EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp . This discovery holds significant implications for our understanding of black hole phenomena.
@@
To determine the trajectory, one must solve geodesic equations with constant radii. The first numerical solutions were obtained by Goldstein goldstein . Mino demonstrated in Mino:2003yg that the geodesic equations can be decoupled, paving the way for Fujita and Hikida to achieve analytical solutions to the timelike geodesic equations in terms of elliptic integrals Fujita:2009bp (see also Battista:2022krl ). Further analytical investigations into Kerr(-Newman) geodesics can be found in works such as Hackmann:2010zz ; Hackmann:2013pva ; Lammerzahl:2015qps ; Wang:2022ouq . Teo leveraged this formalism to derive a one-parameter class of analytical solutions for spherical photon orbits teo and timelike orbits Teo:2020sey . These exact spherical orbits have been extensively explored in subsequent studies (see Grossman:2011ps ; Tavlayan:2020cso ; Tavlayan:2021ylq ; Hod:2013mgr ).
@@
Additionally for null trajectories, a family of parameters exists that causes the photon to return to the same point in the opposite direction after one azimuthal oscillation, satisfying the condition $\Delta\phi=\pi$. This phenomenon, termed the ”photon boomerang,” was introduced by Page Page:2021rhx . In the case of Kerr black holes, it occurs in the near-extremal case, $a\approx 0.994M$, with zero angular momentum. Page demonstrated that the photon boomerang does not occur in naked singularities (NS) for pure Kerr. However, in Anjum:2023axh , it was shown that a Kerr black hole surrounded by perfect fluid dark matter (PFDM) enables the existence of the photon boomerang under NS conditions. Related works on PFDM can be found in Das:2020yxw ; Atamurotov:2021hoq ; Atamurotov:2021hck ; Hou:2018avu ; Haroon:2018ryd ; Das:2021otl ; Narzilloev:2020qtd .
@@
The singularity stands as one of the most generic and conceptually challenging aspects of black holes. However, it is widely believed that the singularity might be resolved or regularized within the framework of quantum gravity. The concept of a regular black hole devoid of a singularity was initially proposed by Sakharov Sakharov:1966aja and later realized by Bardeen Bardeen . In the Bardeen metric, the near-core region exhibits behavior similar to the de Sitter space, and all invariants remain nonsingular throughout. Ayon-Beato and Garcia (ABG) demonstrated that the Bardeen metric can be derived from coupling Einstein’s equations with a nonlinear electrodynamics (NLED) source Ayon-Beato:1998hmi ; Ayon-Beato:2000mjt . For an in-depth review of regular black holes, readers are referred to Ansoldi:2008jw and the accompanying references.
@@
The concept of static regular black holes was subsequently extended to the rotating case by Bambi and Modesto Bambi:2013ufa , as well as by Toshmatov et al. Toshmatov:2014nya using the Newman-Janis algorithm. The resulting metrics are characterized by singularity-free Petrov type D solutions, although they suffer from violations of the weak energy condition. Ghosh Ghosh:2014pba introduced a probability distribution function to mitigate these violations, resulting in a class of three-parameter stationary, symmetric metrics describing regular (nonsingular) rotating black holes. These metrics depend on the mass ($M$), spin ($a$), and a free parameter ($k$) that quantifies the deviation from Kerr and generalizes the Kerr-Newman solution. Extensive literature discussing these solutions and their implications, including constraints from observations by the Event Horizon Telescope (EHT), can be found in Amir:2016cen ; Kumar:2020yem ; Kumar:2020cve ; Kumar:2020ltt ; KumarWalia:2022aop .
@@ II Spherical Orbits around the Kerr-Newman Black Hole
$\begin{split}\Sigma&\equiv r^{2}+a^{2}\cos^{2}{\theta},\\ \Delta&\equiv r^{2}+a^{2}-2Mr+Q^{2},\\ \end{split}$ (3)
and $M$, $Q$, $a$ denote the black hole’s mass, charge, and angular momentum per unit mass, respectively. The allowed range of $a$ $0<a^{2}\leq M^{2}-Q^{2},$ the upper and lower bounds of which are related to the extremal KN and the RN, respectively. The event horizons are $r_{\pm}\equiv M\pm\sqrt{M^{2}-a^{2}-Q^{2}}.$ By defining $u\equiv\cos\theta$ the geodesic equations can be derived using the Hamilton-Jacobi formalism Carter:1968rr ; Johnston:1974pn ; Wang:2022ouq :
@@ II.1.1 Analytical Solutions
the coordinates of $(\theta,\phi,t)$ can be solved in terms of the elliptic integral and the Jacobi elliptic functions Grad Abram Byrd . The Eq. (6) can be cast, by defining $w\equiv u^{2}$, as
@@
and $F(\Psi,k)$ is the incomplete elliptic integral function of the first kind Grad . From Eq. (20) it can be seen that $\lambda$ increases along with $w$. The one full latitudinal oscillation for $\lambda$ is
@@
II.1.2 Non-extreme Kerr-Newman Orbits
The number of KN horizons is parametrized by (II). Here we discuss the non-extremal KN, specifically, we work on the case where
$a=0.8\sqrt{M^{2}-Q^{2}}.$ (33)
The constraints (13)-(14) can be used to determine the orbit parameters. Once both angular momenta for the source,$a$, and the test particle, $\Phi$, are fixed, then the rest of the parameters are completely determined.
In Tables 1-2 we show orbit parameters for $\Phi=0,-M$ and several variations of the charge $Q$.
  Orbit  $Q/M$  $r_{SO}/M$  $C/M^{2}$  $u_{0}$  $\Delta\phi$
  (a)  0  2.671  24.209  1  2.242
  (b)  0.6  2.477  21.546  1  2.050
  (c)   0.7  2.398  20.5271  1  1.941
Table 1: Parameters of the SPO around a non-extreme KN for $\Phi=0$.
In Fig. 2, we present examples of orbit trajectories for two latitudinal oscillations when $\Phi=0$ in the $(x-y-z)$ plane, $x-y$, and $x-z$ planes, respectively. Each plot depicts the orbit on an imaginary sphere with a fixed radius $r_{SO}$. The photons start from the equator and progress northward in each orbit, consistent with the direction of rotation of the black hole itself, which is from west to east. The nine illustrations from Figs. 2 (a), (b), and (c) correspond to cases where $Q=0$ (Kerr), $Q=0.6M$, and $Q=0.7M$, respectively. When $Q=0.6M$ the orbit veers away from the pole. The presence of spherical photon orbits (SPOs) appears to be hindered by the increasing charge. As depicted in Fig. 2 (c), for $Q=0.7M$ the photon orbits predominantly concentrate near the equatorial plane.
Figure 2: Two latitudinal oscillations of the photon orbits around the non-extreme KN with $\Phi=0$ in the $(x-y-z)$ as well as the projective $x-y$ and $x-z$ planes.. The orbit begins from the equator and heads northwards. The direction of the black hole’s rotation is from west to east. Figures (a), (b), and (c) are cases with $Q=0$, $Q=0.6M$, and $Q=0.7M$, respectively.
In Fig. 3 we illustrate three examples of orbits for $\Phi=-M$ in sixteen latitudinal oscillations. Each orbit is plotted on an imaginary sphere with a fixed radius $r_{SO}$ (see Table 2).
  Orbit   $Q/M$  $r_{SO}/M$  $C/M^{2}$   $u_{0}$    $\Delta\phi$
(a)  0  2.882  26.611  0.9821   -0.1922
(b)  0.3  2.833  25.854  0.9815   -0.0299
(c)   0.6  2.669  23.45  0.97967     0.309725
Table 2: Parameters of the SPO around a non-extreme KN for $\Phi=-M$.
In Fig. 3, plots (a), (b), and (c) correspond to the cases of $Q=0$ (Kerr), $Q=0.3M$, and $Q=0.6M$, respectively. Similar to the previous scenario, an increase in charge causes the orbital trajectory to align more closely with the equatorial plane. Of particular interest is the behavior observed for $Q=0.6M$: the orbits densely cluster around a segment of solid angle but do not complete a full azimuthal rotation. This phenomenon arises because $Q=0.6M$ is the transition value between retrograde and prograde motions.
Figure 3: Sixteen latitudinal oscillations of the photon orbits around the non-extreme KN with $\Phi=-M$ in the $(x-y-z)$ as well as the projective $x-y$ and $x-z$ planes. The orbit begins from the equator and heads northwards. The direction of the black hole’s rotation is from west to east. Figures (a), (b), and (c) are cases with $Q=0$, $Q=0.3M$, and $Q=0.6M$, respectively.
II.1.3 Extreme Kerr-Newman Orbit
The extremal KN happens when
@@ II.1.3 Extreme Kerr-Newman Orbit
In this work, the orbit parameters for extremal KN are shown in Tables 3-4 for $\Phi=0,-M$, respectively.
@@
Figs. 5 (a)-(c) depict the scenarios where $Q=0$ (Kerr), $Q=0.5M$, and $Q=0.6M$, respectively. The presence of charge alters the direction of the photon, and its augmentation prompts the orbit to return to the equatorial plane. Notably, for $Q=0.6M$ an intriguing phenomenon happens: the photon undergoes a twist in its rotation direction about the $z$ axis. Initially retrograde ($Q<0.6M$), it abruptly transitions to prograde ($Q\geq 0.6M$). The charge can effectively steer photons to be in the direction of the black hole’s rotation, resulting in a ribbon-shaped orbit, as evidenced in Fig. 5 (c).
@@ III Spherical Orbits around Regular Black Hole
In 1968, Bardeen introduced a regular metric that lacked a physical singularity, exhibiting regularity in both the metric and all invariants as $r\rightarrow 0$ Bardeen . However, Bardeen did not explicitly detail the method used to derive this solution in his original work. Later, Ayon-Beato and Garcia provided an explanation, demonstrating that such a solution can be sourced by a nonlinear electrodynamics (NLED) charge Ayon-Beato:1998hmi ; Ayon-Beato:2000mjt . Since then, numerous proposals have emerged regarding nonsingular black holes coupled with NLED. Balart and Vagenas Balart:2014cga , followed by Culetu Culetu:2014lca , employed a distribution function inspired by the shape of the probability density function to regularize the static metric. The first to extend this regularization approach to the case of a rotating metric in Boyer-Lindquist coordinates was Ghosh Ghosh:2014pba , who derived a class of three-parameter axisymmetric metrics:
@@
Ghosh obtained the solution (III) by transforming the non-singular static black holes Balart:2014cga ; Culetu:2014lca to the rotating solutions via the Newman-Janis algorithm Newman:1965tw . At the core ($r\rightarrow 0$) all the invariants of this metric are regular everywhere. Specifically, the Kretschmann scalar ($K$) approaches zero there, $K(r\rightarrow 0)\rightarrow 0$. For $k=0$ the metric reduces to Kerr (when $a\neq 0$) or Schwarzschild (when $a=0$). In the non-rotating limit ($a\rightarrow 0$) the metric reduces to the regular spherically-symmetric black hole Balart:2014cga ; Culetu:2014lca

### v6

@@ Abstract
We conduct a comprehensive study on spherical orbits around two types of black holes: Kerr-Newman black holes, which are charged, and Ghosh black holes, which are nonsingular. In this work, we consider both null and timelike cases of orbits. Utilizing the Mino formalism, all analytical solutions for the geodesics governing these orbits can be obtained. It turns out that all spherical photon orbits outside the black hole horizons are unstable. In the extremal cases of both models, we obtain the photon boomerangs. The existence of charge in the Kerr-Newman allows the orbits to transition between retrograde and prograde motions, and its increase tends to force the orbits to be more equatorial. On the other hand, the Ghosh black hole, characterized by a regular core and a lack of horizons in certain conditions, presents the possibility of observable stable spherical orbits in the so-called no-horizon condition. As the Ghosh parameter $k$ increases, trajectories tend to exhibit larger latitudinal oscillation amplitudes. We observe that as the Ghosh parameter $k$ increases the trajectories tend to have larger latitudinal oscillation amplitudes. Finally, we investigate the existence of innermost stable spherical orbits (ISSOs). Both black holes demonstrate the appearance of two branches of ISSO radii as a function of the Carter constant $\mathcal{C}$. However, there are notable differences in their behavior: in the case of the Kerr-Newman black hole, the branches merge at a critical value, beyond which no ISSO exists, while for the Ghosh black hole, the transcendental nature of the metric function causes the branches to become complex at some finite distance.
@@ I Introduction
It is well-known that any system possessing spherical symmetry and staticity will preserve its angular momentum, and as a result, the corresponding dynamics are confined on a plane. This is true in classical mechanics as well as in general relativity (GR). What is genuine in GR, among other things, is that not only massive particles but light can also orbit around a massive object. Perhaps the ultimate massive object in the universe is a black hole. Its visual image was first released in 2019 by the Event Horizon Telescope (EHT) collaboration EventHorizonTelescope:2019dse for M87*, and for Sgr A* three years later EventHorizonTelescope:2022wkp . For Schwarzschild we know that the photon circular orbit (photon sphere, $r_{ps}$) is $r_{ps}=3M$, while for Reissner-Nordstrom (RN) there are $r_{ps}^{\pm}=3/2\left(M\pm\sqrt{M^{2}-8Q^{2}/9}\right)$. Both the Schwarzschild and RN unstable photon spheres are outside their corresponding shadow radii, thus can be observed. In reality, however, all astrophysical objects including black holes rotate. The appropriate metric solutions are the Kerr and the Kerr-Newman (KN) Kerr:1963ud ; Newman:1965my . EHT observed the Sgr A* angular diameter $d_{sh}=48.7\pm 7\mu as$, which exhibits consistency with theoretical predictions of Kerr black hole EventHorizonTelescope:2022wkp . Moreover, the observation confirms that the Sgr A* shows the viability of the Kerr model with constrained charge (KN) Ghosh:2022kit ; Vagnozzi:2022moj ; Uniyal:2022vdu
@@
The exploration of ”spherical photon orbits” (SPO) began with Stuchlik’s investigation of the Kerr model Stuchlik1981 , followed by Calvani, De Felice, and Nobili’s study of the Kerr-Newman scenario Calvani:1980is . However, all studies on SPOs suggest that orbits are stable only within the inner horizon. Nonetheless, SPOs play a crucial role in the observation of black holes. It is the multiple unstable photon orbits around black holes that contribute to the formation of the luminous ring detected by instruments like the Event Horizon Telescope (EHT) EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp . This discovery holds significant implications for our understanding of black hole phenomena.
@@
To determine the trajectory, one must solve geodesic equations. The first numerical solutions were obtained by Goldstein goldstein . Mino demonstrated in Mino:2003yg that the geodesic equations can be decoupled, paving the way for Fujita and Hikida to achieve analytical solutions to the timelike geodesic equations in terms of elliptic integrals Fujita:2009bp (see also Battista:2022krl ). Further analytical investigations into Kerr(-Newman) geodesics can be found in works such as Hackmann:2010zz ; Hackmann:2013pva ; Lammerzahl:2015qps ; Wang:2022ouq . Teo leveraged this formalism to derive a one-parameter class of analytical solutions for spherical photon orbits teo and timelike orbits Teo:2020sey . These exact spherical orbits have been extensively explored in subsequent studies (see Grossman:2011ps ; Tavlayan:2020cso ; Tavlayan:2021ylq ; Hod:2013mgr ).
@@
Additionally for null trajectories, a family of parameters exists that causes the photon to return to the same point in the opposite direction after one azimuthal oscillation, satisfying the condition $\Delta\phi=\pi$. This phenomenon, termed the ”photon boomerang,” was introduced by Page Page:2021rhx . In the case of Kerr black holes, it occurs in the near-extremal case, $a\approx 0.994M$, with zero angular momentum. Page demonstrated that the photon boomerang does not occur in naked singularities (NS) for pure Kerr. However, in Anjum:2023axh , it was shown that a Kerr black hole surrounded by perfect fluid dark matter (PFDM) enables the existence of the photon boomerang under NS conditions. Related works on PFDM can be found in Das:2020yxw ; Atamurotov:2021hoq ; Atamurotov:2021hck ; Das:2021otl ; Narzilloev:2020qtd .
@@
Singularity stands as the most generic aspects of black holes. It is, however, widely believed that the singularity might be resolved within the framework of quantum gravity. The concept of a non-singular black hole was initially proposed by Sakharov Sakharov:1966aja and later realized by Bardeen Bardeen . In the Bardeen metric, the near-core region exhibits behavior similar to the de Sitter space, and all invariants remain nonsingular throughout. Ayon-Beato and Garcia (ABG) demonstrated that the Bardeen metric can be derived from coupling Einstein’s equations with a nonlinear electrodynamics (NLED) source Ayon-Beato:1998hmi ; Ayon-Beato:2000mjt . For an in-depth review of regular black holes, readers are referred to Ansoldi:2008jw and the accompanying references.
@@
Regular black holes were subsequently extended to the rotating case by Bambi and Modesto Bambi:2013ufa , as well as by Toshmatov et al. Toshmatov:2014nya using the Newman-Janis algorithm. The resulting metrics are characterized by singularity-free Petrov type D solutions, although they suffer from violations of the weak energy condition. Ghosh Ghosh:2014pba introduced a probability distribution function to mitigate these violations, resulting in a class of three-parameter stationary, symmetric metrics describing regular (nonsingular) rotating black holes. These metrics depend on the mass ($M$), spin ($a$), and a free parameter ($k$) that quantifies the deviation from Kerr and generalizes the Kerr-Newman solution. Extensive literature discussing these solutions and their implications, including constraints from observations by the Event Horizon Telescope (EHT), can be found in Amir:2016cen ; Kumar:2020yem ; Kumar:2020cve ; Kumar:2020ltt .
@@ II Spherical Orbits around the Kerr-Newman Black Hole
$\Sigma$$\equiv$ $r^{2}+a^{2}\cos^{2}{\theta},$
$\Delta$$\equiv$ $r^{2}+a^{2}-2Mr+Q^{2},$ (3)
and $M$, $Q$, $a$ denote the black hole’s mass, charge, and angular momentum per unit mass, respectively. The allowed range of $a$ is
$0\leq a^{2}\leq M^{2}-Q^{2},$ (5)
where the upper and lower bounds correspond to the extremal Kerr-Newman and Reissner-Nordström (RN) black holes, respectively. The event horizons are $r_{\pm}\equiv M\pm\sqrt{M^{2}-a^{2}-Q^{2}}.$ By defining $u\equiv\cos\theta$ the geodesic equations can be derived using the Hamilton-Jacobi formalism Carter:1968rr ; Johnston:1974pn ; Wang:2022ouq :
@@ II.1.1 Analytical Solutions
the coordinates of $(\theta,\phi,t)$ can be solved in terms of the elliptic integral and the Jacobi elliptic functions Grad . The Eq. (8) can be cast, by defining $w\equiv u^{2}$, as
@@
and $F(\Psi,k)$ is the incomplete elliptic integral function of the first kind. From Eq. (22) it can be seen that $\lambda$ increases along with $w$. The one full latitudinal oscillation for $\lambda$ is
@@
II.1.2 Selected Orbits
In this subsection, we focus on the extremal Kerr-Newman (KN) black hole, as the behavior of orbits in the non-extremal case exhibits similar trends to those in the extremal case. The extremal KN black hole occurs when
@@ II.1.2 Selected Orbits
The constraints (15)-(16) can be used to determine the orbit parameters. Once both the angular momentum of the source, $a$, and the angular momentum of the test particle, $\Phi$, are fixed, the remaining parameters are completely determined.
The orbit parameters for the extremal Kerr-Newman (KN) black hole are provided in Tables 1 and 2 for $\Phi=0$ and $\Phi=-M$, respectively.
@@
Figs. 3 (a)-(c) depict the scenarios where $Q=0$ (Kerr), $Q=0.5M$, and $Q=0.6M$, respectively. The presence of charge alters the direction of the photon, and its augmentation prompts the orbit to return to the equatorial plane. Notably, for $Q=0.6M$ an intriguing phenomenon happens: the photon undergoes a twist in its rotation direction about the $z$ axis. Initially retrograde ($Q<0.6M$), it abruptly transitions to prograde ($Q\geq 0.6M$). The charge can effectively steer photons in the direction of the black hole’s rotation, resulting in a ribbon-shaped orbit, as evidenced in Fig. 3 (c).
@@ III Spherical Orbits around Regular Black Hole
In 1968, Bardeen introduced a regular metric that lacked a physical singularity, exhibiting regularity in both the metric and all invariants as $r\rightarrow 0$ Bardeen . However, Bardeen did not explicitly detail the method used to derive this solution in his original work. Later, Ayon-Beato and Garcia provided an explanation, demonstrating that such a solution can be sourced by a nonlinear electrodynamics (NLED) charge Ayon-Beato:1998hmi ; Ayon-Beato:2000mjt . Since then, numerous proposals have emerged regarding nonsingular black holes coupled with NLED. The first to extend this regularization approach to the case of a rotating metric in Boyer-Lindquist coordinates was Ghosh Ghosh:2014pba , who derived a class of three-parameter axisymmetric metrics:
@@
Ghosh obtained the solution (III) by transforming the non-singular static black holes by the shape of the probability density function Balart:2014cga ; Culetu:2014lca to the rotating solutions via the Newman-Janis algorithm Newman:1965tw . At the core ($r\rightarrow 0$) all the invariants of this metric are regular everywhere. Specifically, the Kretschmann scalar ($K$) approaches zero there, $K(r\rightarrow 0)\rightarrow 0$. For $k=0$ the metric reduces to Kerr (when $a\neq 0$) or Schwarzschild (when $a=0$). In the non-rotating limit ($a\rightarrow 0$) the metric reduces to the regular spherically-symmetric black hole Balart:2014cga ; Culetu:2014lca